An experienced mechanic earns $8.00 per hour more than an apprentice. If it takes the apprentice 18 hours to earn the same amount of money as the mechanic can earn in 12 hours, what is the hourly pay rate of the mechanic?

Answer by ewatrrr(24785)

hi,
Let x be the amount apprentice is paid/hr. then (x + 8.00) will be the amount the mechanic is paid/hr
.
Question states:
x/hr*18hr = (x+8)/hr*12hr
.
simplify and solve
18x = 12(x+8)
18x = 12x + 96
6x = 96
x = $16/hr is the apprentice wage
(x+8) = $24/hr the mechanic's wage
.
check your answer
16*18 = 24*12 = 288